At the instant a race began, a 72-kg sprinter exerted a force of 690 N on the starting block at a 22∘ angle with respect to the ground.

A)What was the horizontal acceleration of the sprinter?

B)If the force was exerted for 0.32 s, with what speed did the sprinter leave the starting block?
Answer #1

Horizontal component of force, Fx = 690 cos(22) = 639.76 N

Acceleration, a = Fx/m = 639.76/72 = 8.89 m/s^2

A)

Impulse, I = Ft = 639.76 x 0.32 = 204.72 Ns

B)

Velocity, v = 204.72/72 = 2.84 m/s

Comment in case any doubt please rate my answer....

​​​​​​..